OpenStudy (anonymous):

m= 1/2,b = -1/2 write equation in slop intercept form of the line givin slop m and y intercept b m= 1/2,b = -1/2 write equation in slop intercept form of the line givin slop m and y intercept b @Mathematics

OpenStudy (eyust707):

well do you remember what slope int form is?

OpenStudy (anonymous):

it does nt say

OpenStudy (anonymous):

y=m(x)+b

OpenStudy (anonymous):

\[\le\]

OpenStudy (eyust707):

yes that is the slope intercept for of a line y=mx + b

OpenStudy (eyust707):

so using that plug in your m and b

OpenStudy (eyust707):

what do you get when u do that diamond?

OpenStudy (eyust707):

basically they give you what m equals and the give you what b equals

OpenStudy (eyust707):

it is your job to put them into this equation y=mx + b

OpenStudy (eyust707):

it really just as easy as substituting the letter for a number

OpenStudy (eyust707):

give it a shot diamond

OpenStudy (eyust707):

y= ( ? )x + ( ? )

OpenStudy (anonymous):

\[\ge\]

OpenStudy (eyust707):

im not sure what you mean lol

OpenStudy (eyust707):

the objective of this question is to replace the question marks with the values they give for the m and b